OpenStudy (anonymous):

Find -3a2 - b3 + 3c2 - 2b3 if a = 2, b = -1, and c = 3.

OpenStudy (anonymous):

\(-3a^2 - b^3 + 3c^2 - 2b^3\) \(-3(2)^2 + 1^3 + 3(3)^2 - 2(-1)^3\) Can you simplify that?

OpenStudy (anonymous):

Possibly.

OpenStudy (anonymous):

Okay, first we do the exponents. \(\sf(2)^2, (3)^2~and~(-1)^3\)

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

2^2 = ? 3^2 = ? -1^3 = ? Simplify those

OpenStudy (anonymous):

I don't know that.

OpenStudy (anonymous):

\(2^2 \rightarrow 2 \times 2\) \(3^2 \rightarrow 3 \times 3\) \(-1^3 \rightarrow -1 \times -1 \times -1\)

OpenStudy (secret-ninja):

-3(2)^2 + 1^3 + 3(3)^2 - 2(-1)^3 Use PEMDAS. Parentheses: -3(2) = -6 3(3) = 9 2(-1) = -2 Put them in. -6^2 + 1^3 + 9^2 + 2^3

OpenStudy (secret-ninja):

Now Exponents. -6^2 + 1^3 + 9^2 + 2^3 -6^2 = 36 1^3 = 1 9^2 = 81 2^3 = 8 Put them in. 36 + 1 + 81 + 8

OpenStudy (secret-ninja):

36 + 1 + 81 + 8 Now just add. 36 + 1 = 37 81 + 8 = 89 37 + 89 = 126

OpenStudy (secret-ninja):

Does that help?